# Demystifying Misconceptions in Social Bots Research

### v3

@@ 2 Methodological issues
Annotations of accounts as humans or bots, or any other category, can introduce further bias when the annotation process is opaque or incoherent. As a practical example, cyborgs—accounts that are partly automated and partly managed by humans who can step in as needed to avoid detection—can be labeled as “humans” if one only considers their capacity to reply to a message. However, many bot detectors are based on the definition that any account with at least partial automation is a bot. When evaluating such bot detectors, labeling cyborgs as humans would unfairly inflate the false-positive rate due to the mismatch between the definition adopted by the bot detector versus that of the evaluator. To avoid this bias, account labeling should be performed by multiple independent annotators following a shared and openly accessible rubric, and based on a consistent definition of social bot. Moreover, continuous scores representing the different degrees of “botness” (i.e., automation) should be favored in place of binary labels, as the former are better suited to capture the nuances in the use of automation exhibited by the multitude of online accounts.
Unfortunately, some works in literature fail to follow these principles, resulting in inconsistent labeling and misleading results. Without disclosing any annotation rubric, Gallwitz and Kreil (2022) change criteria for labeling bots multiple times, thus inflating the false-positive or false-negative rates of bot detection methods to fit their narrative. In one case, criticizing Yang et al. (2019), they label accounts that automatically post news headlines—such as those associated with major newswire agencies—as human. Their criticism of Keller and Klinger (2019) goes further. First, Gallwitz and Kreil considered a different set of accounts than those analyzed by Keller and Klinger. Second, they manually inspected only 0.04% of this different set. Third, they ignored suspended accounts, and fourth, they labeled as human accounts those that cross-posted tweets through software apps. The aforementioned four arbitrary choices imply that they did not reproduce Keller and Klinger’s analysis. As such, their conclusion that «not a single one of the accounts in [Keller and Klinger’s] list was a “bot” in any meaningful sense of the word» is simply false. Similarly biased annotations are used by Gallwitz and Kreil (2022) to criticize Dunn et al. (2020), as suspended accounts are again assumed to be human to boost the alleged false-positive rate. Then, accounts that are clearly automated are still presented as classification errors by labeling them as “(unmalicious) bot” and accounts “spreading vaccine-critical information.” The use of obsolete or biased datasets with unclear, inconsistent, or shifting definitions are among the malpractices that affect this field.
@@
The above sections highlighted a number of challenges and widespread methodological issues that currently affect the field of social bots research. Table LABEL:tab:checklist summarizes the main methodological issues and proposes a set of guidelines and practical recommendations to mitigate said issues so as to produce more reliable and robust research. The main issues and recommendations are also discussed in the following.
@@
One significant challenge is the use of contrasting definitions of social bots, which leads to incomparable results and confusion in research. To mitigate this, future studies should openly and clearly define the specific characteristics of the social bots they are targeting. Moreover, future research should adopt definitions consistently, without changing bot criteria and definitions case-by-case. Furthermore, favoring definitions based on objective characteristics—such as the use of automation—over subjective ones, will boost clarity and consistency. When discussing results, researchers should also compare their definitions with existing ones to highlight similarities and differences, also exploring how different definitions could have impacted the results. A related pressing issue is the use of inconsistent or opaque bot labeling procedures, as these hinder the clarity, comparability, and reproducibility of results. Researchers should document and publish their labeling procedures, including the criteria and tools used to carry out the annotation. Sharing labeled datasets with comprehensive metadata will further enhance transparency, reproducibility, and collaborative efforts.
@@
A further set of recommendations involve the evaluation procedures of social bot detectors. Newly trained bot detectors are often tested on well-known and old bots, which does not accurately reflect their performance in real-world scenarios. Researchers should prioritize using the most up-to-date datasets for their evaluations, as these are likely to better reflect the evolving nature of social bots. Performance assessments should include tests on bots with different and unknown characteristics, to better gauge the detectors’ real-world applicability and to ensure that detectors are robust and effective in dynamic online environments. In addition, attention to validation should not only be the focus of the training and testing phases of a new bot detector, but should also be devoted when using a pre-trained detector developed by others, even if it is well-known and established. A subset of automatically assigned labels and scores should always be manually validated, rather than accepted at face value. This ensures their accuracy and reliability in the given context, which might be different than that on which the detector was developed. As with the aforementioned manual labeling procedures, also this manual validation process should be clearly and openly documented, detailing the criteria and methods used for manual verification of the assigned labels.
@@
Another issue in the evaluation of bot detection performance is the possibility to cherry-pick competitors in such a way to purportedly demonstrate the superior performance of a newly proposed detector. To mitigate this issue, future research should benchmark bot detectors against a comprehensive set of existing detectors, including both strong and weak performers. Moreover, the criteria for the selection of the comparisons should be stated clearly and openly. Sharing evaluation scripts and datasets used in benchmarking will further promote transparency and reproducibility in performance comparisons. In addition, sensitivity analyses should also become a standard practice to address the possible sensitivity of detectors to thresholds and other model parameters, which are often overlooked when reporting bot detection results. These analyses should be reported alongside the main results to provide a clearer picture of the robustness and reliability of the detectors, aiding in the identification of optimal settings and limitations.
@@
Lastly, data availability and documentation are crucial for the progress of bot detection research. The lack of shared and well-documented datasets impedes reproducibility and comparison, and slows down the development of new detectors. Researchers should make their datasets publicly available whenever possible, ensuring they are well-documented with comprehensive metadata, including how the data was collected and labelled. This recommendation does not only involve the datasets used to train or evaluate a detector, such as those labeled manually, but also the datasets on which pre-trained detectors are applied. For the latter, it is crucial to share the data together with the labels or scores automatically assigned by the detectors. If data sharing is constrained by privacy or other issues, researchers should at least provide detailed descriptions of their datasets and the procedures for obtaining them. This practice will support reproducibility and facilitate more rigorous comparisons and advancements in bot detection research.
@@
(a) Simplistic and naïve model of social bots and human-operated accounts. Here, all bots are alike, they are evenly distributed, and we have complete knowledge on their numbers and characteristics. According to this model, collecting unbiased and comprehensive bot datasets is possible.
@@ 3 Conceptual issues
Beyond methodological challenges, the landscape of social bots research is also intricately intertwined with a variety of potentially more subtle—yet profoundly influential—conceptual issues. These nuanced challenges, if left unaddressed, have the potential to exert an even more insidious influence on the development of the field and its reception.
@@
In discussing the limitations that currently hinder progress in bot detection, some recent research has pointed the finger at flawed data collection practices that fall short of capturing the complexity of the bot space (Hays et al., 2023). This interpretation of the limitations of the existing datasets implicitly assumes the possibility to encode the full complexity of the bot detection problem space in datasets acquired with rigorous practices, contrary to the literature so far. More specifically, this criticism of the existing data collection practices subsumes that building unbiased and comprehensive bot datasets is possible, and perhaps even convenient. Unbiased datasets are needed so as to avoid that peculiar bot and human characteristics leak into the data allowing simple models to achieve good performance on the collected data, while showing poor performance in real scenarios. Furthermore, comprehensive datasets are needed so that expressive models can learn to generalize to all types of existing social bots. Unfortunately, the assumption of the existence of an unbiased and comprehensive dataset is fallacious.
@@
Regarding bias, many bot datasets were found to contain biases, in that the accounts therein featured some peculiar characteristics (Hays et al., 2023). We note that this is largely unsurprising since such peculiar characteristics are the inevitable consequence of the bots being organized in botnets: groups of accounts created and controlled by a single entity who collectively operates the bots to reach some predefined goal (Zhang et al., 2016). Given that all bots belonging to the same botnet are created and/or operated by the same entity, and pursue the same common goal, they tend to share some similarities. When compared to other accounts, the botnet thus appears to have some peculiar characteristics, as highlighted in Figure 1(b). Thus, in most cases, the presence of biases in the existing bot datasets should not be blamed on the creators of the datasets, but rather should be traced back to the very nature of the social bot phenomenon.
@@
The assumption about the possibility to build comprehensive bot datasets is equally flawed. Creating such datasets would require a uniform and random sample of an adequate number of accounts from the complete distribution of existing bots. This would allow obtaining an accurate representation of the bot landscape that captures the full complexity of the problem space. While in some domains of application for machine learning this is feasible, [^2] For example, in certain speech recognition tasks all instances (e.g., languages) to be classified are known in advance. Given a sufficiently large and diverse set of user inputs, unbiased and comprehensive datasets are possible (Ardila et al., 2020). we argue that performing a uniform and random sample of the whole population of bots is extremely problematic—if not outright impossible. In addition to the many general challenges brought by social data (Olteanu et al., 2019; Pasquetto et al., 2020), the task of detecting social bots also partly belongs to the fields of information security and open-source intelligence, which are intrinsically characterized by the presence of adversaries who are strongly motivated to remain hidden (Cresci et al., 2021). Indeed, there has always been limited knowledge of the real extent of the bot problem, which makes it hard to track bots on a platform, quantify their numbers, and assess their impact (Varol, 2023; Mendoza et al., 2020; Tan et al., 2023). We only have a partial understanding and visibility of the botnets that operate on our online platforms. [^3] The court dispute between Elon Musk and former Twitter CEO Parag Agrawal on the real number of bots on Twitter is a paramount example of the intrinsic difficulties and ambiguity of tracking social bots. Figure 1 provides a conceptual view of the landscape of social bots and human-operated accounts: firstly according to a naïve interpretation (Figure 1(a)) and then in a more realistic representation (Figure 1(b)). Figure 1(b) highlights the knowledge gaps related to the presence of bots in online platforms. These include known unknowns: the exact number of bots belonging to known botnets. But they also include unknown unknowns: information about the hidden botnets operating on a platform. How can we create unbiased and comprehensive datasets when we largely ignore the types and numbers of bots populating online platforms? This fundamental question cannot currently be answered.
The previous discussion highlights the challenges in obtaining a reliable ground-truth for bot detection, which is largely due to the impossibility of acquiring accurate and dependable bot labels. Researchers are often compelled to use proxies, such as platform account removals or manual annotations based on profile characteristics, which however reduce the reliability and validity of the ground-truth. Furthermore, the lack of comprehensive knowledge about the types, sizes, and characteristics of bots operating on platforms creates an imbalance where data about certain bot types may be more accessible than others. Data availability can facilitate research on specific sub-types of bots, potentially leading to more favorable results, albeit confined to a limited subset of bots. However, this could also introduce bias in the development of bot detectors, making them particularly effective against the overrepresented bot types while being ineffective against the underrepresented ones.
@@
Some recent work has interpreted the results of social bot detection studies as a success story for machine learning, concluding that the positive results seem to suggest that the bot detection task is now effectively solved (Hays et al., 2023). Proponents of this perspective highlight the perceived accuracy and reliability of current detection methods, and advocate their suitability for integration into various downstream applications, highlighting the progress made in achieving practical and deployable solutions in the field.
@@
We disagree with this interpretation of the state of social bot research. On the contrary, we believe that unbiased and responsible research in this area should acknowledge the body of literature warning researchers and practitioners about the limitations of bot detection. We are well aware of this body of work, having actively contributed to its development for many years. Below we report on some notable examples. The first accounts of the difficulties supervised detectors face in detecting sophisticated bots date back to 2013 (Yang et al., 2013) and intensified in subsequent years (Cresci et al., 2017; Grimme et al., 2018; Cresci et al., 2019b). The limitations of benchmark datasets for bot detection were discussed by Feng et al. (2021) and the data-related limitations to the generalizability and replicability in social bots research were touched upon by Assenmacher et al. (2022). Several works measured and discussed the generalization deficiencies of bot detectors, with particular emphasis on supervised ones (Echeverría et al., 2018; Yang et al., 2019; Cresci, 2020; Sayyadiharikandeh et al., 2020; Yang et al., 2020; Feng et al., 2021). Others warned about the lack of consistency in bot classifications obtained at different points in time (Rauchfleisch and Kaiser, 2020) and with different detectors (Svenaeus, 2020; Martini et al., 2021). Concerns about the accuracy and usefulness of bot detection have also been raised by security and integrity experts at major platforms (Roth and Pickles, 2020), as well as by independent researchers and the media. [^4] https://www.nytimes.com/2020/06/16/science/social-media-bots-kazemi.html [^5] https://mediawell.ssrc.org/expert-reflections/on-digital-disinformation-and-democratic-myths/ [^6] https://www.lawfareblog.com/random-toxicity-whats-going-benjaminwittess-mentions The growing tension between the continuing challenges of bot detection and the increasing technological capabilities of online manipulators, including unprecedented affordances such as those offered by generative AI (Yang and Menczer, 2023; Ferrara, 2023), even led some researchers to question the long-term viability of the bot detection task (Boneh et al., 2019; Grimme et al., 2022). For the same reasons, some have suggested redirecting part of the ongoing scientific effort away from bot detection and to alternative and more promising tasks, such as detecting information operations and coordinated malicious behavior (Cresci, 2020; Roth and Pickles, 2020). This sampling of the existing literature on the challenges and limitations of bot detection paints a very different picture than some recent, optimistic works. This discrepancy introduces a common misconception about social bots research:
@@
The above evidence shows that despite significant efforts devoted for a prolonged time (Cresci, 2020), bot detection is nowhere near to being a solved problem—quite the contrary.
@@
For many years, we have witnessed a game of whack-a-mole between bot developers, with their increasingly sophisticated accounts (Yang and Menczer, 2023; Yang et al., 2024), and bot hunters, equipped with a variety of different detectors (Yang et al., 2019; Sayyadiharikandeh et al., 2020; Yang et al., 2020). Looking back at how this arms race has unfolded (Cresci et al., 2021), we can conclude that none of the technological advances we have experimented with have significantly mitigated the challenges posed by malicious social bots in the long term. It is reasonable to assume that future advances will suffer a similar fate. Social bots are adversarial, fast-moving targets, characterized by a fast adoption of cutting-edge technology. Bot detection is therefore an intrinsically challenging task, made even more daunting by the lack of accurate information on the targets of the analysis, the limited collaboration from online platforms, and the rapidly evolving nature of online harms.
Below we discuss some other misconceptions that undermine our understanding of social bots and of the related literature:
@@
A common assumption in the scientific literature, as well as among the general public, is that if a given bot detector has achieved good detection performance against some bots, then that detector should be able to replicate that performance on any bot detection task. This assumption is problematic given that there are a plethora of diverse bots, each with its own characteristics (Mazza et al., 2022) and only considering the known botnets. Consider, for example, the differences between the bots used to boost the popularity of certain public figures—so-called fake followers (Cresci et al., 2015)—and those that manipulate trending topics—astroturfers and spammers (Abokhodair et al., 2015). Or the differences between bots involved in political manipulation (Shao et al., 2018; Caldarelli et al., 2020) versus those aimed at fooling automated trading systems (Tardelli et al., 2022; Bello et al., 2023). In other words, not all bots are the same. Some scholars addressed this heterogeneity by designing detectors that aim for generality and broad applicability, such as Botometer, especially in its latest releases (Sayyadiharikandeh et al., 2020; Yang et al., 2020). Others have developed detectors specifically designed to detect certain types of bots. The latter trade generalizability and portability for detection accuracy. Their performance depends heavily on the characteristics of the bots they are designed to detect. For example, a detector designed to detect time-synchronized retweeting actions (Mazza et al., 2019) would likely be useless at detecting mass-following bots (Cresci et al., 2015). However, this should not be regarded as a limitation of the bot detector—let alone one deliberately unstated by its developers in order to pass it off as a good product—but rather as an inappropriate use of the detector itself. Limitations in generalizability and portability also affect general-purpose bot detectors, although to a lesser extent than specialized detectors. In fact, even general-purpose detectors rely on a limited number of features to estimate whether an account is a bot or not. Therefore, in general, any detector, specialized or otherwise, has variable performance and its detection capabilities depend on the characteristics of the accounts it is applied to. In conclusion, no single bot detector is capable of detecting all types of bots.
Many articles proposing a new bot detector, or investigating bot interference, motivate the work on the basis of the role of social bots in spreading disinformation (Shao et al., 2018). Similar arguments are made for the role of bots in exacerbating social conflicts (Stella et al., 2018). In other words, part of the social bot literature is based on the assumption that bots are responsible for many of the problems that affect our online platforms and the information ecosystem as a whole. This brings to the fore another misconception:
@@
An unbiased analysis of the existing literature paints a much more dubious picture of the role of social bots. For example, while some studies have concluded that bots play a prominent role in the spread of problematic content (Shao et al., 2018; Stella et al., 2018), others have reached the opposite conclusion (Vosoughi et al., 2018; González-Bailón and De Domenico, 2021; Seckin et al., 2024). The existing literature has focused almost exclusively on detecting bots and characterizing their behavior, leaving the fundamental task of measuring the impact of bot malfeasance largely unexplored (Cresci, 2020). For these reasons, we currently lack scientific consensus and conclusive evidence about the role of social bots and their effectiveness in influencing online users. What we do know, however, is that bots are only one of many agents involved in the spread of misinformation and disinformation (Starbird, 2019; Roth and Pickles, 2020). Examples of some of the other agents include state-sponsored trolls, users who collude and coordinate for malicious purposes, superspreaders, and even willing but unwitting individuals (Starbird, 2019; DeVerna et al., 2022). Each of these agents represents a potential threat to safe and trusted online platforms, and a thriving area of research and experimentation. It is therefore critical to balance efforts in all of these directions, avoiding the pitfall of overusing some while overlooking others, based on unsupported decisions.
The previous misconceptions might induce readers to think that social bot research has led to flawed—if not outright useless—results. Such a claim has been made by Hays et al. (2023). Similarly, Gallwitz and Kreil (2022) not only question the validity of existing bot detection methods—based on the fallacies discussed earlier—but go even further in arguing that social bots are not an issue. These claims lead to our last, but by no means least important, misconception:
@@
reference main result validation or corroboration source
Yang et al. (2014) studied social bots and developed effective detection techniques platform collaborations in real-world scenarios platform deployments in production environments
Subrahmanian et al. (2016) hosted and presented results of the DARPA Twitter bot challenge findings incorporated into defense and cybersecurity strategies U.S. Department of Defense reports and cybersecurity policies
Ferrara et al. (2016) brought to light the issue of social bots and their capabilities in manipulation confirmed from various data leaks showing coordinated bot operations federal investigation reports, media disclosures
Cresci et al. (2017) demonstrated that social bots evolve, becoming more sophisticated with time supported by subsequent studies and data showing similar patterns academic follow-up studies and industry research
Varol et al. (2017) estimated social bots prevalence reports and admissions by social media platforms public statements and transparency reports from social media
Lazer et al. (2018) highlighted the impact of bots on misinformation spread and public opinion alignment with findings from investigations into electoral interference government investigation documents and hearings
Badawy et al. (2019) analyzed use of bots in Russian interference with the 2016 U.S. election corroborated by U.S. intelligence community findings U.S. Senate and House intelligence committee reports
Cresci (2020) framed social bot detection as an adversarial machine learning task demonstrated by the improvements obtained via adversarial approaches academic follow-up studies and industry research
Ferrara (2022) provided a taxonomy of Twitter bots, methods, and tools to detect them results supported by platforms and public organizations reports and investigations by government agencies
Table 2: Examples of key findings on social bots that have withstood the test of time. Each result is accompanied by details on its validation or corroboration, along with the supporting sources. While not exhaustive, this list underscores the impact that rigorous social bots research can achieve.

### v4

@@ 2 Methodological issues
Annotations of accounts as humans or bots, or any other category, can introduce further bias when the annotation process is opaque or incoherent. As a practical example, cyborgs—accounts that are partly automated and partly managed by humans who can step in as needed to avoid detection—can be labeled as “humans” if one only considers their capacity to reply to a message. However, many bot detectors are based on the definition that any account with partial automation is a bot. When evaluating such bot detectors, labeling cyborgs as humans would unfairly inflate the false-positive rate due to the mismatch between the definition adopted by the bot detector versus that of the evaluator. In other cases, the authors change criteria for labeling bots multiple times within the context of the same study, without disclosing any annotation rubric (Gallwitz and Kreil, 2022). First, they label accounts that automatically post news headlines—such as those associated with major newswire agencies—as humans. Then, they ignore suspended accounts or consider them to be human, depending on the analysis. Finally, they label accounts that cross-post tweets through software apps as humans.
As exemplified above, the use of obsolete or biased datasets with unclear, inconsistent, or shifting definitions and opaque labeling schemes are among the malpractices that affect this field. To avoid bias, account labeling should be performed by multiple independent annotators following a shared and openly accessible rubric, and based on a consistent definition of social bot. Moreover, continuous scores representing the different degrees of “botness” (i.e., automation) should be favored in place of binary labels, as the former are better suited to capture nuances in the use of automation.
@@
Table LABEL:tab:checklist summarizes the main methodological issues that may occur at each step of the social bots research process, and proposes a set of best practices, guidelines, and practical recommendations to mitigate them.
@@
One significant challenge is the use of contrasting definitions of social bots, which leads to incomparable results and confusion in research. To mitigate this, future studies should openly and clearly define the specific characteristics of the social bots they are targeting. Moreover, future research should adopt definitions consistently, without changing bot criteria and definitions case-by-case. Furthermore, favoring definitions based on objective characteristics—such as the use of automation—over subjective ones, will boost clarity and consistency. When discussing results, researchers should also compare their definitions with existing ones to highlight similarities and differences, also exploring how different definitions could impact the results. A related pressing issue is the use of inconsistent or opaque bot labeling procedures, as these hinder the clarity, comparability, and reproducibility of results. Researchers should document and publish their labeling procedures, including the criteria and tools used to carry out the annotation. Sharing labeled datasets with comprehensive metadata will further enhance transparency, reproducibility, and collaborative efforts.
@@
A further set of recommendations involves the evaluation procedures of social bot detectors. Newly trained bot detectors are often tested on well-known and old bots, which does not accurately reflect their performance in real-world scenarios. Researchers should prioritize using the most up-to-date datasets for their evaluations, as these are likely to better reflect the evolving nature of social bots. Performance assessments should include tests on bots with different and unknown characteristics, to better gauge the detectors’ real-world applicability and to ensure that detectors are robust and effective in dynamic online environments. In addition, attention to validation should not only be the focus of the training and testing phases of a new bot detector, but should also be devoted when using a pre-trained detector developed by others, even if it is well-known and established. A subset of automatically assigned labels and scores should always be manually validated, rather than accepted at face value. This ensures their accuracy and reliability in the given context, which might be different than that on which the detector was developed. As with the aforementioned manual labeling procedures, also this manual validation process should be clearly and openly documented, detailing the criteria and methods used for manual verification of the assigned labels.
@@
Another issue in the evaluation of bot detection performance is the possibility of cherry-picking competitors in such a way to purportedly demonstrate the superior performance of a newly proposed detector. To mitigate this issue, future research should benchmark bot detectors against a comprehensive set of existing detectors, including both strong and weak performers. Moreover, the criteria for the selection of the comparisons should be stated clearly and openly. Sharing evaluation scripts and datasets used in benchmarking will further promote transparency and reproducibility in performance comparisons. In addition, sensitivity analyses should also become a standard practice to address the possible sensitivity of detectors to thresholds and other model parameters, which are often overlooked when reporting bot detection results. These analyses should be reported alongside the main results to provide a clearer picture of the robustness and reliability of the detectors, aiding in the identification of optimal settings and limitations.
@@
Lastly, data availability and documentation are crucial for the progress of bot detection research. The lack of shared and well-documented datasets impedes reproducibility and comparison, and slows down the development of new detectors. Researchers should make their datasets publicly available whenever possible, ensuring they are well-documented with comprehensive metadata, including how the data was collected and labeled. This recommendation does not only involve the datasets used to train or evaluate a detector, such as those labeled manually, but also the datasets on which pre-trained detectors are applied. For the latter, it is crucial to share the data together with the labels or scores automatically assigned by the detectors. If data sharing is constrained by privacy or other issues, researchers should at least provide detailed descriptions of their datasets and the procedures for obtaining them. This practice will support reproducibility and facilitate more rigorous comparisons and advancements in bot detection research.
@@
(a) Simplistic and naïve model of social bots (black-colored) and human-operated accounts (grey-colored). Here, all bots are alike, they are evenly distributed, and we have complete knowledge of their numbers and characteristics. According to this model, collecting unbiased and comprehensive bot datasets is possible.
@@ 3 Conceptual issues
Beyond methodological challenges, the field is also intricately intertwined with a variety of potentially more subtle—yet profoundly influential—conceptual issues. These concern how social bots and related phenomena are defined, framed, and understood within the research discourse. These nuanced challenges, if left unaddressed, have the potential to exert an even more insidious influence on the development of the field and its reception.
@@
In discussing the limitations that currently hinder progress in bot detection, some recent research pointed at flawed data collection practices that fall short of capturing the complexity of the bot space (Hays et al., 2023). This interpretation of the limitations of the existing datasets implicitly assumes the possibility to encode the full complexity of the bot detection problem space in a dataset. More specifically, this criticism subsumes that building unbiased and comprehensive bot datasets is possible, and perhaps even convenient. Unbiased datasets are needed so as to avoid that peculiar bot and human characteristics leak into the data allowing simple models to achieve good performance on the collected data, while showing poor performance in real scenarios. Furthermore, comprehensive datasets are needed so that expressive models can learn to generalize to all types of existing social bots. Unfortunately, the assumption of the existence of an unbiased and comprehensive dataset is fallacious.
@@
Regarding bias, many bot datasets were found to contain biases, in that the accounts therein featured some peculiar characteristics (Hays et al., 2023). However, this is largely unsurprising since such characteristics are the inevitable consequence of the bots being organized in botnets: groups of accounts created and controlled by a central entity who collectively operates the bots to reach some predefined goal (Zhang et al., 2016). Given that all bots belonging to the same botnet are created or operated by the same entity, and pursue the same common goal, they tend to share similarities. When compared to other accounts, the botnet thus appears to have some peculiar characteristics, as highlighted in Figure 1(b). Thus, in most cases, the presence of biases in the existing bot datasets should not be blamed on the creators of the datasets, but rather traced back to the very nature of the social bot phenomenon.
@@
The assumption about the possibility to build comprehensive bot datasets is equally flawed. Creating such datasets would require a uniform and random sample of an adequate number of accounts from the complete distribution of existing bots. This would allow obtaining an accurate representation of the bot landscape that captures the full complexity of the problem space. However, performing a uniform and random sample of the whole population of bots is extremely problematic—if not outright impossible. The task of detecting social bots partly belongs to the fields of information security and open-source intelligence, which are intrinsically characterized by the presence of adversaries who are strongly motivated to remain hidden (Cresci et al., 2021). There has always been limited knowledge of the real extent of the bot problem, which makes it hard to track bots on a platform, quantify their numbers, and assess their impact (Varol, 2023; Mendoza et al., 2020; Tan et al., 2023). We only have a partial understanding and visibility of the botnets that operate on our online platforms. [^2] The court dispute between Elon Musk and former Twitter CEO Parag Agrawal on the real number of bots on Twitter is a paramount example of the intrinsic difficulties and ambiguity of tracking social bots. Figure 1 provides a conceptual view of the landscape of social bots and human-operated accounts: firstly according to a naïve interpretation (Figure 1(a)) and then in a more realistic representation (Figure 1(b)). Figure 1(b) highlights the knowledge gaps related to the presence of bots in online platforms. These include known unknowns: the exact number of bots belonging to known botnets. But they also include unknown unknowns: information about the hidden botnets operating on a platform. How can we create unbiased and comprehensive datasets when we largely ignore the types and numbers of bots populating online platforms? This fundamental question is currently unanswered.
@@
Some recent work posed the idea that the seemingly excellent results in social bot detection could be interpreted as a success story for machine learning and as evidence that the task is now effectively solved (Hays et al., 2023). According to this perspective, the perceived accuracy and reliability of current detection methods support their suitability for integration into various downstream applications, highlighting the progress made in achieving practical and deployable solutions. While Hays et al. (Hays et al., 2023) pose the question “Is bot detection a solved problem?” as a rhetorical device, both their framing and conclusions imply that the field has collectively overlooked critical challenges. This is a problematic foundation. A vast literature has long acknowledged the limitations in social bots research. By sidestepping this body of work, the authors construct a straw man: a field that blindly celebrates near-perfect performance without reflection. Other than boosting the apparent novelty of their critique, this framing misrepresents the field. Unbiased and responsible research in this area should acknowledge the body of literature warning researchers and practitioners about the limitations of bot detection. Below, we report on some notable examples. The first accounts of the difficulties supervised detectors face in detecting sophisticated bots date back to 2013 (Yang et al., 2013) and intensified in subsequent years (Cresci et al., 2017; Grimme et al., 2018; Cresci et al., 2019b). The limitations of benchmark datasets for bot detection were discussed by Feng et al. (Feng et al., 2021), and the data-related limitations to the generalizability and replicability in social bots research were touched upon by Assenmacher et al. (Assenmacher et al., 2022). Several works measured and discussed the generalization deficiencies of bot detectors, with particular emphasis on supervised ones (Echeverr´ıa et al., 2018; Yang et al., 2019; Cresci, 2020; Sayyadiharikandeh et al., 2020; Yang et al., 2020; Feng et al., 2021). Others warned about the lack of consistency in bot classifications obtained at different points in time (Rauchfleisch and Kaiser, 2020) and with different detectors (Svenaeus, 2020; Martini et al., 2021). Concerns about the accuracy and usefulness of bot detection have also been raised by security and integrity experts at major platforms (Roth and Pickles, 2020), and by independent researchers and the media. [^3] https://www.nytimes.com/2020/06/16/science/social-media-bots-kazemi.html [^4] https://mediawell.ssrc.org/expert-reflections/on-digital-disinformation-and-democratic-myths/ [^5] https://www.lawfareblog.com/random-toxicity-whats-going-benjaminwittess-mentions The growing tension between the lasting challenges of bot detection and the increasing technological capabilities of online manipulators, including unprecedented affordances such as those offered by generative AI (Yang and Menczer, 2023; Ferrara, 2023), even led some researchers to question the long-term viability of the bot detection task (Boneh et al., 2019; Grimme et al., 2022). For the same reasons, some have suggested redirecting part of the ongoing scientific effort away from bot detection and to alternative and more promising tasks, such as detecting information operations and coordinated harmful behavior (Cresci, 2020; Roth and Pickles, 2020; Mannocci et al., 2024). This sampling of the existing literature on the challenges and limitations of bot detection paints a very different picture than some recent, optimistic works. This discrepancy introduces a common misconception about social bots research:
@@
The above evidence shows that despite significant efforts devoted for a prolonged time, bot detection is nowhere near to being a solved problem—quite the contrary.
@@
For many years, we witnessed to a whack-a-mole game between bot developers, with their increasingly sophisticated accounts (Yang and Menczer, 2023; Yang et al., 2024), and bot hunters, equipped with a variety of different detectors (Yang et al., 2019; Sayyadiharikandeh et al., 2020; Yang et al., 2020). Looking back at how this arms race has unfolded, we can conclude that none of the technological advances we have experimented with have significantly mitigated the challenges posed by malicious social bots in the long term. It is reasonable to assume that future advances will suffer a similar fate. Social bots are adversarial, fast-moving targets, characterized by a fast adoption of cutting-edge technology (Cresci et al., 2021). Bot detection is therefore an intrinsically challenging task, made even more daunting by the lack of accurate information on the targets of the analysis, the limited collaboration from online platforms, and the rapidly evolving nature of online harms.
Below, we discuss some other misconceptions that undermine the understanding of social bots and of the related literature:
@@
Both scholars and the general public often assume that a bot detector performing well on some bots will perform equally well on any bot detection task. This assumption is problematic given that there are a plethora of diverse bots, each with its own characteristics (Mazza et al., 2022). Consider, for example, the differences between the bots used to boost the popularity of certain public figures—so-called fake followers (Cresci et al., 2015)—and those that manipulate trending topics—astroturfers and spammers (Abokhodair et al., 2015). Or the differences between bots involved in political manipulation (Shao et al., 2018; Caldarelli et al., 2020) versus those aimed at fooling automated trading systems (Tardelli et al., 2022; Bello et al., 2023). Some scholars addressed this heterogeneity by designing detectors that aim for generality and broad applicability, such as Botometer, especially in its latest releases (Sayyadiharikandeh et al., 2020; Yang et al., 2020). Others have developed detectors specifically designed to detect certain types of bots. The latter trade generalizability and portability for detection accuracy, and their performance depends heavily on the characteristics of the bots they are designed to detect. For example, a detector designed to detect time-synchronized retweeting actions (Mazza et al., 2019) would likely be useless at detecting mass-following bots (Cresci et al., 2015). However, this should not be regarded as a limitation of the bot detector—let alone one deliberately unstated by its developers in order to pass it off as a good product—but rather as an inappropriate use of the detector itself. Limitations in generalizability and portability also affect general-purpose bot detectors, although to a lesser extent than specialized ones. In fact, even general-purpose detectors rely on a limited number of features to estimate whether an account is a bot. Therefore, in general, any detector, specialized or otherwise, has variable performance and its detection capabilities depend on the characteristics of the accounts it is applied to. In conclusion, no single bot detector is capable of detecting all types of bots.
Social bots research is often framed within the broader landscape of mis- and disinformation studies, a field that extends well beyond online manipulation, including critical perspectives on post-truth dynamics, cognitive biases, epistemic trust, ideological polarization, and the structural features of media ecosystems (Spohr, 2017; Waisbord, 2018). The specific relevance of social bots in this broader discourse has largely emerged from concerns over their potential to enable scalable and low-cost manipulation (Shao et al., 2018; Stella et al., 2018). These concerns, while widely shared and influential, are themselves being increasingly reexamined, raising questions about the actual impact of bots, the assumptions behind their perceived threat, and the relative emphasis they have received in disinformation research. This surfaces another misconception:
@@
An unbiased analysis of the existing literature paints a dubious picture of the role of social bots. For example, while some studies have concluded that bots play a prominent role in the spread of problematic content (Shao et al., 2018; Stella et al., 2018), others have reached the opposite conclusion (Vosoughi et al., 2018; González-Bailón and De Domenico, 2021; Seckin et al., 2024). The existing literature has focused almost exclusively on detecting bots and characterizing their behavior, leaving the fundamental task of measuring the impact of bot malfeasance largely unexplored (Cresci, 2020). For these reasons, we currently lack scientific consensus and conclusive evidence about the role of social bots and their effectiveness in influencing online users. What we do know, however, is that bots are only one of many agents involved in the spread of mis- and disinformation (Starbird, 2019; Roth and Pickles, 2020). Examples of other agents are state-sponsored trolls, users who collude and coordinate for malicious purposes, superspreaders, and even willing but unwitting individuals (Starbird, 2019; DeVerna et al., 2022). Each of these agents represents a potential threat to safe and trusted online platforms, and a thriving area of research and experimentation. It is therefore critical to balance efforts in all of these directions, avoiding the pitfall of overstudying some while overlooking others, based on unsupported decisions.
The previous misconceptions might induce readers to think that social bot research has led to flawed—if not outright useless—results. Such claims have been recently made by Hays et al. (Hays et al., 2023) and Gallwitz & Kreil (Gallwitz and Kreil, 2022)—the latter even in sensationalist terms. These claims lead to our last, but by no means least important, misconception: